OpenStudy (surryyy):

If we toss a coin 100 times, then we expect that heads will appear 50 times. Answer True or False

OpenStudy (anonymous):

Trueeeeeeee

OpenStudy (surryyy):

r u sure?

OpenStudy (anonymous):

Pr (Head) = 1/2 Pr (Tail) = 1/2 Half of 100 is 50, therefore TRUE

OpenStudy (anonymous):

yes it is true 100 C 50

OpenStudy (surryyy):

ok thank u guys!